f(x)=3x-2 and g(x)=1/(2x+3)
solve for (f*g)(-2)

Answer by nik_consult(24)

f(x)=3x-2 and g(x)=1/(2x+3)
solve for (f*g)(-2)
g(-2) = [1/ 2(-2)+3]
= [1/-1]
= -1
therefore fg(-2)= f[g(-2)]
= f[-1]
= 3(-1)-2
=-3-2
= -5
therefore fg(-2)= -5